# Learning Robust and Correct Controllers Guided by Feasibility-Aware Signal Temporal Logic via BarrierNet

### v2

@@ III Problem Formulation and Approach
Existing works [nguyen2016exponential, xiao2021high] employ HOCBFs (7) to address safety specifications in (8d) with high relative degree by embedding them as constraints in quadratic optimization-based controllers. In these approaches, HOCBF constraints are enforced as strict hard constraints to guarantee safety, which implicitly requires the QP to remain feasible at every sampling instant. However, this feasibility is not guaranteed and is often violated when the control bound $\mathcal{U}$ are tight or when the HOCBF hyperparameters (e.g., the coefficients in the class $\kappa$ functions) are manually designed and fixed. In particular, based on (5), fixed hyperparameters force the constraint $\psi_{i}(\mathbf{x},t)\geq 0$ to impose a constant decay rate on function $\psi_{i-1}(\mathbf{x},t)$, regardless of how close the system actually is to the boundary. This leads to overly aggressive constraint satisfaction in non-critical situations, resulting in unnecessarily conservative control actions and reduced performance. In [xiao2021barriernet], BarrierNet was introduced as a differentiable QP layer with HOCBF constraints appended to the output of a neural network controller to guarantee safety. In this framework, class $\kappa$ hyperparameters are learned during training, which substantially reduces conservativeness compared with fixed designs. However, the approach relies on supervised learning, where a reference control input that already satisfies safety and input bounds must be provided during training. Such a reference input is often difficult to obtain, and becomes especially challenging when the task is governed by complex STL specifications (8d). Therefore, there is a strong need for methods that integrate BarrierNet with STL satisfaction while automatically generating reference inputs, rather than requiring them to be specified a priori.
@@
Approach: In this paper, we train a neural network controller, with or without memory as in (3) and (2), to solve the optimization problem in Problem 1. We address STL satisfaction by enforcing trainable HOCBF‐based constraints during both training and execution. From the STL formula $\varphi$, we construct time‐varying HOCBFs and embed them into a neural network controller via a BarrierNet layer [xiao2021barriernet], enabling specification correctness and robustness to be incorporated directly into policy learning. The HOCBF hyperparameters are generated by three networks: InitNet, which adapts them to different initial conditions; RefNet, which automatically produces reference‐input updates along the rollout; and extended BarrierNet architecture, inspired by [xiao2021barriernet], which adjusts CBF hyperparameters to enforce $\varphi$. These components are jointly optimized using a unified robustness metric that accounts for both STL satisfaction and QP feasibility, allowing the resulting controller to remain feasible even under tight input bounds or complex specifications. After training, the networks provide hyperparameters online, ensuring feasible STL enforcement without manual tuning or pre‐generated reference inputs, thereby reducing conservativeness and improving robustness in both training and execution.
IV Joint Learning of Feasibility-Aware Signal Temporal Logic and BarrierNet
@@
In this section, we present our solution to Problem 1. Section IV-A introduces the trainable HOCBFs and the state-dependent BarrierNet architecture extended from [19], along with a sufficient condition for ensuring QP feasibility. Section IV-B provides a general procedure for constructing a set of time-varying HOCBFs that guarantee the satisfaction of a given STL specification. Section IV-C then defines a unified robustness metric that captures both STL satisfaction and QP feasibility. Finally, Section IV-D describes how these time-varying HOCBFs are jointly trained with the neural network controller through BarrierNet to further enhance robustness.
@@ IV-A Extended BarrierNet with Trainable HOCBFs and Feasibility Guarantees
where $\alpha_{i,j}$ are class $\kappa$ functions, $\infty>p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})>0$ for $i=1,\ldots,m$ and $j=1,\ldots,M$, and $m$ denotes the relative degree of the HOCBF $b_{j}$. For notational simplicity, we write $\psi_{\cdot,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})$ to denote $\psi_{\cdot,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p},\mathbf{x}_{0},\mathbf{P}_{\text{inip}})$ throughout the remainder of this paper. Each $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})$ is time-varying, as it depends on the state $\mathbf{x}(t)$, while its initial value $\mathbf{P}_{\text{inip}}(\mathbf{x}_{0},\boldsymbol{\theta}_{\text{inip}})$ depends on the initial condition and trainable parameter $\boldsymbol{\theta}_{\text{inip}}$ (This will also be clarified in Section IV-D).
BarrierNet [xiao2021barriernet] is a neural network layer implemented as a differentiable Quadratic Program (dQP) with HOCBF constraints. We append an extended BarrierNet layer, extending the formulation in [xiao2021barriernet], as the terminal layer of the neural network controller (with or without memory), as shown in (3) and (2), where $\mathbf{u}^{*}(t)$ is obtained from:
@@
where $\mathbf{Q}(\cdot,\boldsymbol{\theta}_{q})\in\mathbb{R}^{q\times q}$, $\mathbf{F}(\cdot,\boldsymbol{\theta}_{f})\in\mathbb{R}^{q}$, $\infty>p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})>0$ for $i=1,\ldots,m$, and $b_{j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\mathbf{x}_{0})$ are all generated by preceding neural network layers with trainable parameters $(\boldsymbol{\theta}_{q},\boldsymbol{\theta}_{f},\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{\text{inip}},\boldsymbol{\theta}_{p})\coloneqq\boldsymbol{\theta}$. Here, the symbol $\cdot$” indicates the input to the network: $\mathbf{x}(t)$ in the memoryless case, and the state history $\mathbf{x}_{0:t}$ when memory is used. The matrix $\mathbf{Q}$ is positive definite, and $\mathbf{Q}^{-1}\mathbf{F}$ can be interpreted as a reference control. Although $\mathbf{Q}$, $\mathbf{F}$, and $p_{i,j}$ are generated by previous layers in (10), they may also be treated as directly trainable parameters. The dQP (10) is solved at each time point $k\Delta t$, $k=0,1,\ldots$ until reaching the time horizon $T$, and the solution $\mathbf{u}^{*}(t)$ is applied to the system as a constant for the time period $k\Delta t,k\Delta t+\Delta t)$.
@@
Given an time-varying HOCBF $b(\mathbf{x},t)$ with a sequence of sets $\mathcal{C}_{1}(t),\ldots,C_{m}(t)$ as defined in (6), if $\mathbf{x}(0)\in\mathcal{C}_{1}(0)\cap\mathcal{C}_{2}(0)\cap\ldots\cap\mathcal{C}_{m}(0)$, and $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})$, $i\in\{1,\ldots,m-2\}$, are differentiable and chosen such that all their time derivatives along system (1) that appear in the recursion up to $\psi_{m-1,j}$ are independent of the control input $\mathbf{u}$, i.e.,
$L_{g}L_{f}^{k}p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})=0,k=m-2-i,$ (11)
$\forall i=1,\ldots,m-2,m\geq 3,$
so that Lipschitz controller $\mathbf{u}$ can enter the HOCBF constraint only through the highest-order terms in $\psi_{m,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})$ (10) under (10c), or if $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})$ are treated as trainable hyperparameters that do not depend on the state, then a BarrierNet composed of neurons as in (10) guarantees the forward invariance of the intersection of sets in (6).
@@
We show that each constraint (10) has the standard HOCBF form, thus forward invariance follows from Thm. 2.
Assume that $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})$, $i\in\{1,\dots,m-2\}$, are differentiable and satisfy (11) (the multiplier $p_{m-1,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})$ can introduce a dependence on $\mathbf{u}$ only in $\psi_{m,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})$), or are trainable hyperparameters. By repeatedly applying the modified recursion in (9), the highest–order constraint $\psi_{m,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})$ associated with $b_{j}(\mathbf{x},t,\bm{\theta}_{b},\mathbf{x}_{0})$ takes the form
@@
Since (10) is differentiable, the gradients of the optimal control $\mathbf{u}^{*}(t)$ and the HOCBF hyperparameters $p_{i,j}$ with respect to $\boldsymbol{\theta}$ can be obtained via the dQP framework of [amos2017optnet] (the gradients of $p_{i,j}$ with respect to $\mathbf{x}$ can also be computed by PyTorch). This enables $\boldsymbol{\theta}$ to be trained with any common neural–network optimizer (e.g., those used for LSTMs or CNNs [goodfellow2016deep]). As a result, BarrierNet not only guarantees satisfaction of all HOCBF constraints encoding the STL requirements, but also enables the controller to optimize a given objective through training.
@@
Given an time-varying HOCBF $b(\mathbf{x},t)$ with a sequence of sets $\mathcal{C}_{1}(t),\ldots,C_{m}(t)$ as defined in (6), and $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})$, $i\in\{1,\ldots,m-2\}$, are differentiable and satisfy (11), or just trainable hyperparameters, assume $\mathbf{x}(0)\in\mathcal{C}_{1}(0)\cap\mathcal{C}_{2}(0)\cap\ldots\cap\mathcal{C}_{m}(0)$. For each $j\in\{1,\dots,M\}$, let the highest-order HOCBF constraint be defined by (10) and $\mathbf{u}$ satisfies (10c). If $\psi_{m-1,j}>0$ is enforced for all $t\geq 0$, then the optimization problem in (10) is pointwise feasible for any $(\mathbf{x},t)$ with $\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})>0$, implying that the dQP admits a feasible control $\mathbf{u}^{*}(t)$.
@@
The condition $\psi_{m-1,j}>0$ in Thm. 3 is a sufficient condition for the feasibility of problem (10), and, together with Thm. 2, guarantees the forward invariance of the HOCBF-defined set and feasibility of the QP at every time step. From (18), we observe that if $\psi_{m-1,j}>0$ cannot be maintained, one can instead impose $\psi_{m-2,j}>0$ as the feasibility condition. In this case, by increasing $L_{f}p_{m-1,j}$, the left-hand side of (18) can still be rendered positive, ensuring the feasibility of the constraint. If $\psi_{m-2,j}>0$ also cannot be ensured, we may proceed hierarchically and use $\psi_{m-3,j}>0$ as the feasibility condition, and so on, down to $\psi_{0,j}>0$.
@@
Different from the original BarrierNet [xiao2021barriernet], which only considers time-invariant HOCBFs $b_{j}(x)$, the formulation in (10) incorporates time-varying HOCBFs $b_{j}(x,t)$. Moreover, we make the HOCBF $b_{j}$ itself trainable, in addition to the hyperparameters $\mathbf{Q}$, $\mathbf{F}$, and $p_{i,j}$. In Theorem 2, we explicitly allow the time derivatives of the multipliers $p_{i,j}$ to introduce dependence on the input $\mathbf{u}$ in the highest–order constraint $\psi_{m,j}$, a feature that was not considered in [xiao2021barriernet]. Therefore, the BarrierNet structure in (10), together with Theorem 2, extends the original BarrierNet framework proposed in [xiao2021barriernet]. In [liu2023learning], the trainable multiplier $p_{i,j}(\mathbf{x}_{0},\boldsymbol{\theta}_{p})$ remains constant over time because it depends only on the initial state, although it varies across rollouts. This limits its ability to adapt along the time horizon and restricts potential robustness improvement. In contrast, in this paper $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{init}})$ is time-varying, allowing it to continuously adjust to the evolving system state and thereby better satisfy the STL specification.

### v3

@@ III Problem Formulation and Approach
STL formulas provide a unified and expressive framework to specify both safety requirements (e.g., obstacle avoidance) and goal-reaching tasks (e.g., reach within a specified time interval) over system trajectories. Existing works [28, 38] employ HOCBFs (7) to address safety specifications in (8d) with high relative degree by embedding them as constraints in quadratic optimization-based controllers. In these approaches, HOCBF constraints are enforced as strict hard constraints to guarantee safety, which implicitly requires the QP to remain feasible at every sampling instant. However, feasibility is not guaranteed and is often violated when the control bounds $\mathcal{U}$ are tight or when the HOCBF hyperparameters (e.g., the coefficients in the class $\kappa$ functions) are manually designed and fixed. In particular, based on (5), fixed hyperparameters force the constraint $\psi_{i}(\mathbf{x},t)\geq 0$ to impose a constant decay rate on function $\psi_{i-1}(\mathbf{x},t)$, regardless of how close the system actually is to the boundary. This leads to overly aggressive constraint satisfaction in non-critical situations, resulting in unnecessarily conservative control actions and reduced performance. In [39], BarrierNet was introduced as a differentiable QP layer with HOCBF constraints appended to the output of a neural network controller to guarantee safety. In this framework, class $\kappa$ hyperparameters are learned during training, which substantially reduces conservativeness compared with fixed designs. However, the approach relies on supervised learning, where a reference control input that already satisfies safety and input bounds must be provided during training. Such a reference input is often difficult to obtain, and becomes especially challenging when the task is governed by complex STL specifications. Therefore, there is a strong need for methods that integrate BarrierNet with STL satisfaction while automatically generating reference inputs, rather than requiring them to be specified a priori.
@@
Approach: In this paper, we train a neural network controller, with or without memory as in (3) and (2), to solve the optimization Problem 1. We address STL satisfaction by enforcing trainable HOCBF‐based constraints during both training and execution. From the STL formula $\varphi$, we construct time‐varying HOCBFs and embed them into a neural network controller via a BarrierNet layer [39], enabling specification correctness and robustness to be incorporated directly into policy learning. The HOCBF hyperparameters are generated by three networks: InitNet, which adapts them to different initial conditions; RefNet, which automatically produces reference controls; and extended BarrierNet, inspired by [39], which adjusts CBF hyperparameters to enforce $\varphi$. These components are jointly optimized using a unified robustness metric that accounts for both STL satisfaction and QP feasibility, allowing the resulting controller to remain feasible even under tight input bounds or complex specifications. After training, the networks provide hyperparameters online, ensuring feasible STL enforcement without manual tuning or pre‐generated reference inputs, thereby reducing conservativeness and improving robustness in both training and execution. An overview of the proposed approach is illustrated in Fig. 1, while additional implementation details are provided in Sec. IV.
IV Learning Controllers from Feasibility-Aware STL via BarrierNet
@@
In this section, we present our solution to Problem 1. Section IV-A introduces the trainable HOCBFs and the state-dependent BarrierNet architecture extended from [19], along with a sufficient condition for guaranteeing QP feasibility. Section IV-B provides a general procedure for constructing a set of time-varying HOCBFs that guarantee the satisfaction of a given STL specification. Section IV-C then defines a unified robustness metric that captures both STL satisfaction and QP feasibility. Finally, Section IV-D describes how these time-varying HOCBFs are jointly trained with the neural network controller through BarrierNet to further enhance robustness.
@@ IV-A Extended BarrierNet with Trainable HOCBFs and Feasibility Guarantees
where $\alpha_{i,j}$ are class $\kappa$ functions, $\infty>p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})>0$ for $i=1,\ldots,m$ and $j=1,\ldots,M$, and $m$ denotes the relative degree of the HOCBF $b_{j}$. For notational simplicity, we write $\psi_{\cdot,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})$ to denote $\psi_{\cdot,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p},\mathbf{x}_{0},\mathbf{P}_{\text{inip}})$ throughout the remainder of this paper. Each $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\mathrm{inip}})$ is time-varying, as it depends on the state $\mathbf{x}(t)$ and the trainable parameter $\boldsymbol{\theta}_{p}$. Its initial value $\mathbf{P}_{\mathrm{inip}}(\mathbf{x}_{0},\boldsymbol{\theta}_{\mathrm{inip}})$ depends on the initial condition and the trainable parameter $\boldsymbol{\theta}_{\mathrm{inip}}$ (This will also be clarified in Section IV-D).
BarrierNet [39] is a neural network layer implemented as a differentiable Quadratic Program (dQP) with HOCBF constraints. We append an extended BarrierNet layer, extending the formulation in [39], as the terminal layer of the neural network controller (with or without memory), as shown in (3) and (2), where $\mathbf{u}^{*}(t)$ is obtained from:
@@
where $\mathbf{Q}(\cdot,\boldsymbol{\theta}_{q})\in\mathbb{R}^{q\times q}$, $\mathbf{F}(\cdot,\boldsymbol{\theta}_{f})\in\mathbb{R}^{q}$, $\infty>p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})>0$ for $i=1,\ldots,m$, and $b_{j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\mathbf{x}_{0})$ are all generated by preceding neural network layers with trainable parameters $(\boldsymbol{\theta}_{q},\boldsymbol{\theta}_{f},\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{\text{inip}},\boldsymbol{\theta}_{p})\coloneqq\boldsymbol{\theta}$. Here, the symbol $\cdot$” indicates the input to the network: $\mathbf{x}(t)$ in the memoryless case, and the state history $\mathbf{x}_{0:t}$ when memory is used. The matrix $\mathbf{Q}$ is positive definite, and $\mathbf{Q}^{-1}\mathbf{F}$ can be interpreted as a reference control. Although $\mathbf{Q}$, $\mathbf{F}$, and $p_{i,j}$ are generated by previous layers in (10), they may also be treated as directly trainable hyperparameters. The dQP (10) is solved at each time point $k\Delta t$, $k=0,1,\ldots$ until reaching the time horizon $T$, and the solution $\mathbf{u}^{*}(t)$ is applied to the system as a constant for the time period $k\Delta t,k\Delta t+\Delta t)$.
@@
Given a time-varying HOCBF $b(\mathbf{x},t)$ with a sequence of sets $\mathcal{C}_{1}(t),\ldots,\mathcal{C}_{m}(t)$ defined in (6), assume that the initial condition satisfies $\mathbf{x}(0)\in\mathcal{C}_{1}(0)\cap\mathcal{C}_{2}(0)\cap\cdots\cap\mathcal{C}_{m}(0).$ Let $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})$, $i\in\{1,\ldots,m-2\}$ be differentiable functions chosen such that all time derivatives appearing in the recursion up to $\psi_{m-1,j}$ are independent of the control input $\mathbf{u}$. Specifically, we require
$L_{g}L_{f}^{k}p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})=0,~k=m-2-i,$ (11)
$m\geq 3,~\forall i=1,\ldots,m-2.$
Under this condition, the Lipschitz continuous control input $\mathbf{u}$ enters the HOCBF constraint only through the highest-order term $\psi_{m,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})$ in (10), subject to the input bound (10c). As a result, forward invariance of the set intersection defined in (6) is guaranteed regardless of whether the functions $p_{i,j}$ depend on the state. In particular, when $p_{i,j}$ are treated as trainable hyperparameters that do not depend on the state, the resulting BarrierNet composed of neurons defined in (10) also guarantees forward invariance.
@@
We show that each constraint (10) has the HOCBF form, thus forward invariance follows from Theorem 2.
Figure 1: Overall structure of the controller. The purple module runs only at $t=0$, while the blue modules operate at every later time step. The dashed box outlines the controller $\pi(\cdot,\boldsymbol{\theta})$ ((2) or (3)) over one rollout (epoch).
Assume that $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})$, $i\in\{1,\dots,m-2\}$ are differentiable and satisfy (11) (the multiplier $p_{m-1,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})$ can introduce a dependence on $\mathbf{u}$ only in $\psi_{m,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})$), or are trainable hyperparameters. By repeatedly applying the modified recursion in (9), the highest–order constraint $\psi_{m,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})$ associated with $b_{j}(\mathbf{x},t,\bm{\theta}_{b},\mathbf{x}_{0})$ takes the form
@@
Since (10) is differentiable, the gradients of the optimal control $\mathbf{u}^{*}(t)$ and the HOCBF hyperparameters $p_{i,j}$ with respect to $\boldsymbol{\theta}$ can be obtained via the dQP framework of [4] (the gradients of $p_{i,j}$ with respect to $\mathbf{x}$ can also be computed by PyTorch). This enables $\boldsymbol{\theta}$ to be trained with any common neural–network optimizer (e.g., those used for LSTMs or CNNs [10]). As a result, BarrierNet not only guarantees satisfaction of all HOCBF constraints encoding the STL requirements, but also enables the controller to optimize a given objective through training.
@@
Consider a time-varying HOCBF $b(\mathbf{x},t)$ and a sequence of sets $\mathcal{C}_{1}(t),\ldots,\mathcal{C}_{m}(t)$ defined in (6). Assume that the initial condition satisfies $\mathbf{x}(0)\in\mathcal{C}_{1}(0)\cap\mathcal{C}_{2}(0)\cap\cdots\cap\mathcal{C}_{m}(0).$ For $i\in\{1,\ldots,m-2\}$, let $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{inip}})$ be differentiable functions that satisfy (11). Alternatively, these functions may be treated as trainable hyperparameters that do not depend on the state. For each $j\in\{1,\ldots,M\}$, define the highest-order HOCBF constraint by (10), and assume that the control input $\mathbf{u}$ satisfies the bound (10c). Suppose that the condition $\psi_{m-1,j}>0$ is enforced for all $t\geq 0$. Then, for any $(\mathbf{x},t)$ such that $\psi_{m-1,j}(\mathbf{x},t,\boldsymbol{\theta}_{b},\boldsymbol{\theta}_{p})>0$, the optimization problem in (10) is pointwise feasible. As a result, the corresponding dQP admits a feasible control input $\mathbf{u}^{*}(t)$.
@@
The condition $\psi_{m-1,j}>0$ in Theorem 3 is a sufficient condition for the feasibility of problem (10), and, together with Theorem 2, guarantees the forward invariance of the HOCBF-defined set and feasibility of the QP at every time step. From (18), we observe that if $\psi_{m-1,j}>0$ cannot be maintained, one can instead impose $\psi_{m-2,j}>0$ as the feasibility condition. In this case, by increasing $L_{f}p_{m-1,j}$, the left-hand side of (18) can still be rendered positive, ensuring the feasibility of the constraint. If $\psi_{m-2,j}>0$ also cannot be ensured, we may proceed hierarchically and use $\psi_{m-3,j}>0$ as the feasibility condition, and so on, down to $\psi_{0,j}>0$, i.e., $b_{j}>0$.
@@
Different from the original BarrierNet [39], which only considers time-invariant HOCBFs $b_{j}(\mathbf{x})$, the formulation in (10) incorporates time-varying HOCBFs $b_{j}(\mathbf{x},t)$. Moreover, we make the HOCBF $b_{j}$ itself trainable, in addition to the hyperparameters $\mathbf{Q}$, $\mathbf{F}$, and $p_{i,j}$. In Theorem 2, we explicitly allow the time derivatives of the multipliers $p_{i,j}$ to introduce dependence on the input $\mathbf{u}$ in the highest–order constraint $\psi_{m,j}$, a feature that was not considered in [39]. Therefore, the BarrierNet structure in (10), together with Theorem 2, extends the original BarrierNet framework proposed in [39]. In [24], the trainable multiplier $p_{i,j}(\mathbf{x}_{0},\boldsymbol{\theta}_{p})$ remains constant over time because it depends only on the initial state, although it varies across rollouts. This limits its ability to adapt along the time horizon and restricts potential robustness improvement. In contrast, in this paper $p_{i,j}(\mathbf{x},\boldsymbol{\theta}_{p},\mathbf{P}_{\text{init}})$ is time-varying, allowing it to continuously adjust to the evolving system state and thereby better satisfy the STL specification.